# Uncovering Quantum Many-body Scars with Quantum Machine Learning

### v1

@@ Abstract
Quantum many-body scars are rare eigenstates hidden within the chaotic spectra of many-body systems, representing a weak violation of the eigenstate thermalization hypothesis (ETH). Identifying these scars, as well as other non-thermal states in complex quantum systems, remains a significant challenge. Besides exact scar states, the nature of other non-thermal states lacking simple analytical characterization remains an open question. In this study, we employ tools from quantum machine learning—specifically, quantum convolutional neural networks (QCNNs), to explore hidden non-thermal states in chaotic many-body systems. Our simulations demonstrate that QCNNs achieve over 99% single-shot measurement accuracy in identifying all known scars. Furthermore, we successfully identify new non-thermal states in models such as the xorX model, the PXP model, and the far-coupling Su-Schrieffer-Heeger model. In the xorX model, some of these non-thermal states can be approximately described as spin-wave modes of specific quasiparticles. We further develop effective tight-binding Hamiltonians within the quasiparticle subspace to capture key features of these many-body eigenstates. Finally, we validate the performance of QCNNs on IBM quantum devices, achieving single-shot measurement accuracy exceeding 63% under real-world noise and errors, with the aid of error mitigation techniques. Our results underscore the potential of QCNNs to uncover hidden non-thermal states in quantum many-body systems.
@@ I Introduction
Chaos is a fascinating phenomenon in both classical and quantum systems. In chaotic systems, small perturbations can lead to vastly different dynamical trajectories due to the extreme sensitivity to initial conditions, making long-term predictions highly challenging. The ergodicity and mixing properties of chaotic dynamics underpin the principles of statistical physics and thermodynamics [1, 2, 3]. Occasionally, amidst the chaotic spectrum, there are sometimes integrable, periodic trajectories known as “scars”, where the system temporarily exhibits regular behavior. In contrast, integrable systems are more straightforward to analyze, as their motions are fully predictable and often display periodic dynamics. Besides fully chaotic and integrable dynamics, there exists a class of pseudo-integrable dynamics, where certain chaotic characteristics are present despite some Lyapunov exponents being zero [4, 5]. These dynamics blend regular and complex behavior, sitting between full integrability and complete chaos.
@@
In certain quantum many-body systems, the eigenstate thermalization hypothesis (ETH) can be weakly violated, allowing non-chaotic eigenstates to persist within an otherwise chaotic spectrum. These rare eigenstates, known as quantum many-body scars (QMBSs), represent a vanishing fraction of the entire eigenstates [12, 13, 14]. In contrast to chaotic eigenstates, QMBSs offer an intriguing window into non-thermal behavior, posing a challenge for traditional analytical methods. Experimental studies with a 51-atom quantum simulator have observed persistent revivals over long timescales [15], attributed to an evenly spaced tower of QMBSs [12, 13]. These states are characterized by sub-volume-law entanglement entropy, setting them apart from the typical chaotic eigenstates [16]. The strict definition of QMBS remains an open problem, complicated by the complex nature of many-body systems. Identifying atypical eigenstates, which could be candidates for QMBSs, is a challenging task. It is pivotal for understanding ergodicity breaking in quantum systems. Some of these states may be linked to pseudo-integrability due to finite energy and Hilbert space [17, 18, 19]. The weak integrability breaking can exhibit longer thermalization times in quantum many body systems [20, 21].
Quantum machine learning (QML) offers powerful tools to study quantum-many-body physics. Among these, quantum convolutional neural networks (QCNNs) have been particularly effective in classifying different phases of matter in quantum spin systems, as demonstrated both theoretically [22, 23] and experimentally [24]. In particular, QCNNs enable model-independent learning, where the phase boundary can be predicted from synthetic fixed-point states associated with each phase individually [25]. However, identifying quantum many-body scars (QMBSs) presents a greater challenge than identifying phases due to the scarcity of data, as QMBSs constitute only a vanishingly small fraction of the eigenstates. Recent studies [26, 27] have applied unsupervised learning to investigate scars, while Ref. [28] has generated scar states by manually engineering conserved quantities. Despite these advances, many other non-thermal states in these systems remain unexplored, highlighting the need for further investigation into hidden eigenstates that could exhibit atypical properties.
@@
Figure 1: (color online) Conceptual plot of the non-thermal states classified by QCNN. The wave functions inside and outside the scar subspace are treated as two classes that are input to the QCNN as training data. The QCNN consists of convolution layers(C), pooling layers(P) and a fully connected layer(F). Then QCNN is able to identify a larger subspace (dashed circles) of special states that are potential candidates for approximate scars.
II Models
In this work, we consider three different systems harboring QMBSs, as detailed in this section. Our primary focus is the xorX model. In this model, single spin flip (X) occurs when its nearest neighbors satisfy the exclusive or (xor) condition. The xorX model stands out because it allows for the analytical solution of a family of exact scar states, providing a clear testbed for exploring non-thermal phenomena, though it remains an open question whether other types of scar states exist. The xorX model under open boundary condition (OBC) is described by the Hamiltonian [16]
$H_{\rm xorX}=\lambda\sum_{i=2}^{n-1}\left(\sigma_{i}^{x}-\sigma_{i-1}^{z}\sigma_{i}^{x}\sigma_{i+1}^{z}\right)+\Delta\sum_{i=1}^{n}\sigma_{i}^{z}+J\sum_{i=1}^{n-1}\sigma_{i}^{z}\sigma_{i+1}^{z}~{},$ (1)
@@ II Models
$|\mathcal{S}_{m}\rangle$ $=$ $\frac{1}{m!\sqrt{\mathcal{N}_{m}}}\left(Q^{\dagger}\right)^{m}|0\rangle^{\otimes n}~{},$ (2)
@@
$Q^{\dagger}$ $=$ $\sum_{i=2}^{n-1}(-1)^{i}P_{i-1}^{0}\sigma_{i}^{+}P_{i+1}^{0}~{},$ (3)
@@
$H_{\rm PXP}=\frac{\Omega}{2}\sum_{i=2}^{n-1}P_{i-1}^{0}\sigma_{i}^{x}P_{i+1}^{0}~{},$ (4)
@@
$+J_{\rm nn}\sum_{i=1}^{n-3}\sigma_{i}^{+}\sigma_{i+3}^{-}+{\rm h.c.}~{}.$
@@
III Enhancement of Quantum Convolutional Neural Network
We further enhance the QCNN to improve its ability to classify quantum states with greater precision. Specifically, we introduce additional convolution layers after each existing convolution layer to enable the network to capture correlations over longer distances, analogous to the effect of using larger kernels in classical convolutional neural networks. The performance with different numbers of convolutional layers is shown in Fig. 2. To tackle more complex classification tasks, we introduce a general layer of universal quantum gates with a constant depth to preprocess the input data. This adjustment helps handle finite-size effect and outliers which breaks translational-symmetry, such as disorders or boundaries. Furthermore, we incorporate an ancillary qubit initialized in the $|0\rangle$ state, analogous to the zero padding at the boundary pixels in classical convolutional neural networks. This allows us to perform classification operations without altering the size of the Hilbert space (see Appendix A).
Figure 2: (color online) Training the xorX model with varying numbers of trainable gate parameters, denoted as $n_{\theta}$. The rate of marked states $r_{m}$ is the ratio of marked states to the total dimension of the Hilbert space. Each point represents a distinct set of circuit parameters and a corresponding training iteration. The orange dashed line indicates the rate of marked (scar) states in the input data. The system consists of 12 spins. The number of trainable parameters is given by $n_{\theta}=12+3(9n_{l}+3)$, where $n_{l}$ is the number of convolutional layers preceding the pooling layers.
IV Training with Exact Scar States
$\mathcal{L}$ $=$ $\frac{1}{d}\sum_{i=0}^{d-1}\left|y_{i}-q_{i}\right|~{},$ (6)
where $d$ is the total number of input states, and $q_{i}$ denotes the probability of the output qubit being in |$\ket{1}$ state. This $q_{i}$ also corresponds to the single-shot measurement accuracy for identifying exact scar states. The enhanced QCNN is expected to distinguish scar states from thermal states after sufficient training with enough number of layers. After training, we evaluate the QCNN’s performance by testing it on the eigenstates of the Hamiltonian. For convenience, we consider an eigenstate to be “marked” if the output $q_{i}$ exceeds 50%, matching the rate of scar states in the training set. Experimentally, this classification is achieved through a majority vote across multiple trials and measurements [38]. In this context, achieving zero loss indicates perfect classification, although zero classification error does not necessarily imply the loss function fully decays to zero. The loss function also reflects the model’s robustness to noise.
@@ IV Training with Exact Scar States
Figure 3: (color online) The entanglement entropy and inverse participation ratio of the eigenstates of xorX model within different domain-wall number sectors. (a)(c) $n_{\rm dw}=2$, (b)(d) $n_{\rm dw}=3$. Other parameters are $\lambda=J=10\Delta$. The red crosses are the eigenstates marked by the QCNN. The prisms are the exact scar states. The number of spins is 12 while two of them at the boundaries are fixed.
@@
We begin our study with the xorX model in Eq. (1), where there is a family of well-defined exact scar states. Since the exact scar states Eq. (2) are independent of the parameters in the Hamiltonian Eq. (1), the trained QCNN is also parameter independent. Before experimental implementation that will be presented in Sec. VI, we first perform numerical simulations on classical computers. After training, the quantum circuit classifies the eigenstates into two types. Interestingly, while the total loss decays during the training, the final converged loss remains near 0.14. Moreover, it successfully recognizes all the exact scar states with an error probability of single-shot measurement less than 1%. This means that the QCNN definitely recognizes all the exact scar state and is expected to do so with sufficient measurement in experiment [38]. The disparity between the large loss function and the high accuracy in recognizing scar states indicates that the QCNN also classifies some additional states, beyond the known exact scars, as “scar” states. As we detail in Sec. V, we identify a substantial portion of these states as non-thermal states, which bear a significant resemblance to the exact scar states.
The additional non-thermal states have similar energy as the exact scar states, as shown in Fig. 3. These states are situated in the middle of the energy spectrum, distinguishing them from the low-energy integrable modes. Their half-chain entanglement entropy is lower than that of the bulk chaotic states, as shown in Fig. 3(a)(b), indicating potential deviation from the volume-law entanglement entropy. The QCNN can compensate for deficiencies that the entanglement entropy may fail to distinguish states [39]. Additionally, the inverse participation ratio (IPR), defined as $\sum_{i}\left|\langle\psi|i\rangle\right|^{4}$ for a given state |$\ket{\psi}$ in the computational basis {|$\{\ket{i}\}$, is significantly higher than that of the majority of chaotic states, as shown in Fig. 3(c)(d). They are thus constrained within a smaller Hilbert space compared to chaotic states [40].
@@
Figure 4: (color online) Fidelity as a function of time of an initial state that is an equal amplitude superposition of exact scar states (blue dashed line), additional states marked by QCNN (red solid line), and non-marked states (green dash-dotted line).
@@
Figure 5: (color online) The proportion of marked eigenstates identified by the QCNN to the total eigenstates. The results are shown for architectures with 2 (blue solid line) and 3 (orange dashed line) convolutional layers before the pooling layers.
@@
V Spin-wave approximation for the marked states in xorX model
The numerical results presented in the previous section suggest a more detailed study on the nature of the additional non-thermal states found by QCNN in the xorX model in Eq. (1). The IPR indicates that these states are predominantly localized within a small subregion of the full Hilbert space. In this section, we demonstrate that some of these non-thermal states can be understood in terms of quasiparticles, specifically magnon bound states. We will construct effective tight-binding Hamiltonians that approximately describe the spin-wave modes of these quasiparticles, reproducing key features of the exact many-body eigenstates.
V.1 Integrable states
Figure 6: (color online) Domain wall dynamics between two antiferromagnetically ordered domains, with the inset highlighting the eigenstates within this subspace. These eigenstates are characterized by distinct eigen-energies with similar entanglement entropy.
@@ V.1 Integrable states
$H=\sum_{i=1}^{n-1}(-1)^{i}\Delta d_{i}^{\dagger}d_{i}+\lambda d_{i+1}^{\dagger}d_{i}+\lambda d_{i}^{\dagger}d_{i+1}~{}.$ (7)
@@
$H$ $=$ $\left(\begin{matrix}\Delta&\lambda+\lambda e^{ik}\\ \lambda+\lambda e^{-ik}&-\Delta\end{matrix}\right)~{},$ (8)
where we have set the lattice spacing to unity. The eigenenergies are $E_{k}=\pm\sqrt{\Delta^{2}+4\lambda^{2}\cos^{2}(k/2)}$, with corresponding eigenstates $\phi_{k}$. Under OBC, the eigenstates approximate standing waves, expressed as superpositions of $\phi_{k}$ and $\phi_{-k}$. Specifically, these superpositions take the form of $\left(\phi_{k}+\phi_{-k}\right)/\sqrt{2}$ and $\left(\phi_{k}-\phi_{-k}\right)/\sqrt{2}$. These states exhibit low entanglement entropy, characteristic of integrable systems, as shown in the inset of Fig. 6. The QCNN successfully identifies states within this integrable subspace, marking those with energies similar to the exact scar states.
Figure 7: (color online) The mean and variance of the total $S_{z}$ for each eigenstate within a certain $n_{\rm dw}$ sector: (a)&(c) $n_{\rm dw}=2$; (b)&(d) $n_{\rm dw}=3$. The red crosses indicate the eigenstates identified by the QCNN. The green solid curves represent analytical results from the ferromagnetic magnon bound state approximation, while the purple dashed curves correspond to those obtained from the anti-ferromagnetic magnon bound state approximation. Other parameters are $\lambda=J=10\Delta$.
V.2 Approximate quasiparticle states

### v2

@@ Abstract
Quantum many-body scars are rare eigenstates hidden within the chaotic spectra of many-body systems, representing a weak violation of the eigenstate thermalization hypothesis (ETH). Identifying these scars, as well as other non-thermal states in complex quantum systems, remains a significant challenge. Besides exact scar states, the nature of other non-thermal states lacking simple analytical characterization remains an open question. In this study, we employ tools from quantum machine learning—specifically, (enhanced) quantum convolutional neural networks (QCNNs), to explore hidden non-thermal states in chaotic many-body systems. Our simulations demonstrate that QCNNs achieve over 99% single-shot measurement accuracy in identifying all known scars. Furthermore, we successfully identify new non-thermal states in models such as the xorX model, the PXP model, and the far-coupling Su-Schrieffer-Heeger model. In the xorX model, some of these non-thermal states can be approximately described as spin-wave modes of specific quasiparticles. We further develop effective tight-binding Hamiltonians within the quasiparticle subspace to capture key features of these many-body eigenstates. Finally, we validate the performance of QCNNs on IBM quantum devices, achieving single-shot measurement accuracy exceeding 63% under real-world noise and errors, with the aid of error mitigation techniques. Our results underscore the potential of QCNNs to uncover hidden non-thermal states in quantum many-body systems.
@@ I INTRODUCTION
Chaos is a fascinating phenomenon in both classical and quantum systems. In chaotic systems, small perturbations can lead to vastly different dynamical trajectories due to the extreme sensitivity to initial conditions, making long-term predictions highly challenging. The ergodicity and mixing properties of chaotic dynamics underpin the principles of statistical physics and thermodynamics [1, 2, 3]. Occasionally, amidst the chaotic spectrum, there are sometimes integrable, periodic trajectories known as “scars", where the system temporarily exhibits regular behavior. In contrast, integrable systems are more straightforward to analyze, as their motions are fully predictable and often display periodic dynamics. Besides fully chaotic and integrable dynamics, there exists a class of pseudo-integrable dynamics, where certain chaotic characteristics are present despite some Lyapunov exponents being zero [4, 5]. These dynamics blend regular and complex behavior, sitting between full integrability and complete chaos.
@@
In certain quantum many-body systems, the eigenstate thermalization hypothesis (ETH) can be weakly violated, allowing non-chaotic eigenstates to persist within an otherwise chaotic spectrum. These rare eigenstates, known as quantum many-body scars (QMBSs), represent a vanishing fraction of the entire eigenstates [12, 13, 14, 15]. In contrast to chaotic eigenstates, QMBSs offer an intriguing window into non-thermal behavior, posing a challenge for traditional analytical methods. Experimental studies with a 51-atom quantum simulator have observed persistent revivals over long timescales [16], attributed to an evenly spaced tower of QMBSs [12, 13]. These states are characterized by sub-volume-law entanglement entropy, setting them apart from the typical chaotic eigenstates [17]. The strict definition of QMBS remains an open problem, complicated by the complex nature of many-body systems. Identifying atypical eigenstates, which could be candidates for QMBSs, is a challenging task. It is pivotal for understanding ergodicity breaking in quantum systems. Some of these states may be linked to pseudo-integrability due to finite energy and Hilbert space [18, 19, 20]. The weak integrability breaking can exhibit longer thermalization times in quantum many body systems [21, 22].
Quantum machine learning (QML) offers powerful tools to study quantum-many-body physics. Among these, quantum convolutional neural networks (QCNNs) are designed to detect patterns in quantum systems and have demonstrated advantages over classical convolutional neural networks when processing classical input data [23]. QCNNs have proven particularly effective in classifying different phases of matter in quantum spin systems, as demonstrated both theoretically [24, 25] and experimentally [26]. In particular, QCNNs enable model-independent learning, where the phase boundary can be predicted from synthetic fixed-point states associated with each phase individually [27]. However, identifying quantum many-body scars (QMBSs) presents a greater challenge than identifying phases due to the scarcity of data, as QMBSs constitute only a vanishingly small fraction of the eigenstates. Recent studies [28, 29] have applied unsupervised learning to investigate scars, while Ref. [30] has generated scar states by manually engineering conserved quantities. Despite these advances, many other non-thermal states in these systems remain unexplored, highlighting the need for further investigation into hidden eigenstates that could exhibit atypical properties.
Figure 1: Conceptual plot of the non-thermal states classified by QCNN. The wave functions inside and outside the scar subspace are treated as two classes that are input to the QCNN as training data. The QCNN consists of convolution layers(C), pooling layers(P) and a fully connected layer(F). Then QCNN is able to identify a larger subspace (dashed circles) of special states that are potential candidates for approximate scars.
@@
II RESULTS
II.1 Models
In this work, we consider three different systems harboring QMBSs, as detailed in this section. Our primary focus is the xorX model. In this model, single spin flip (X) occurs when its nearest neighbors satisfy the exclusive or (xor) condition. The xorX model stands out because it allows for the analytical solution of a family of exact scar states, providing a clear testbed for exploring non-thermal phenomena, though it remains an open question whether other types of scar states exist. The xorX model under open boundary condition (OBC) is described by the Hamiltonian [17, 30]
$H_{\rm xorX}=\lambda\sum_{i=2}^{n-1}\left(\sigma_{i}^{x}-\sigma_{i-1}^{z}\sigma_{i}^{x}\sigma_{i+1}^{z}\right)+\Delta\sum_{i=1}^{n}\sigma_{i}^{z}+J\sum_{i=1}^{n-1}\sigma_{i}^{z}\sigma_{i+1}^{z}\leavevmode\nobreak\ ,$ (1)
@@ II.1 Models
$|\mathcal{S}_{m}\rangle$ $=$ $\frac{1}{m!\sqrt{\mathcal{N}_{m}}}\left(Q^{\dagger}\right)^{m}|0\rangle^{\otimes n}\leavevmode\nobreak\ ,$ (2)
@@
$Q^{\dagger}$ $=$ $\sum_{i=2}^{n-1}(-1)^{i}P_{i-1}^{0}\sigma_{i}^{+}P_{i+1}^{0}\leavevmode\nobreak\ ,$ (3)
@@
$H_{\rm PXP}=\frac{\Omega}{2}\sum_{i=2}^{n-1}P_{i-1}^{0}\sigma_{i}^{x}P_{i+1}^{0}\leavevmode\nobreak\ ,$ (4)
@@
$+J_{\rm nn}\sum_{i=1}^{n-3}\sigma_{i}^{+}\sigma_{i+3}^{-}+{\rm h.c.}\leavevmode\nobreak\ .$
@@
Figure 2: The entanglement entropy and participation ratio of the eigenstates of xorX model within different domain-wall number sectors. (a)(c) $n_{\rm dw}=2$, (b)(d) $n_{\rm dw}=3$. Other parameters are $\lambda=J=10\Delta$. The red crosses are the eigenstates marked by the QCNN. The prisms are the exact scar states. The number of spins is 12 while two of them at the boundaries are fixed.
II.2 Non-thermal states
@@
We begin our study with the xorX model in Eq. (1), where there is a family of well-defined exact scar states. Since the exact scar states Eq. (2) are independent of the parameters in the Hamiltonian Eq. (1), the trained QCNN is also parameter independent. Before experimental implementation that will be presented in section Experimental demonstration on quantum device, we first perform numerical simulations on classical computers. After training, the quantum circuit classifies the eigenstates into two types. Interestingly, while the total loss decays during the training, the final converged loss remains near 0.14. Moreover, it successfully recognizes all the exact scar states with an error probability of single-shot measurement less than 1%. This means that the QCNN definitely recognizes all the exact scar state and is expected to do so with sufficient measurement in experiment [37]. The disparity between the large loss function and the high accuracy in recognizing scar states indicates that the QCNN also classifies some additional states, beyond the known exact scars, as “scar" states. As we detail in section Spin-wave approximation for the marked states in xorX model, we identify a substantial portion of these states as non-thermal states, which bear a significant resemblance to the exact scar states.
@@
Figure 3: Revival behavior under Hamiltonian evolution. The plot shows fidelity as a function of time, with the initial state being an equal-amplitude superposition of exact scar states (blue dashed line), additional states identified by QCNN (red solid line), and non-marked states (green dash-dotted line). The parameters are the same as in Figure 2.
Figure 4: The proportion of marked eigenstates identified by the QCNN to the total eigenstates. The results are shown for architectures with 2 (blue solid line) and 3 (orange dashed line) convolutional layers before the pooling layers.
The additional non-thermal states have similar energy as the exact scar states, as shown in Fig. 2. These states are situated in the middle of the energy spectrum, distinguishing them from the low-energy integrable modes. Their half-chain entanglement entropy is lower than that of the bulk chaotic states, as shown in Fig. 2(a)(b), indicating potential deviation from the volume-law entanglement entropy. The QCNN can compensate for deficiencies that the entanglement entropy may fail to distinguish states [38]. Additionally, the participation ratio (PR), defined as $\sum_{i}\left|\langle\psi|i\rangle\right|^{4}$ for a given state |$\ket{\psi}$ in the computational basis {|$\{\ket{i}\}$, is significantly higher than that of the majority of chaotic states, as shown in Fig. 2(c)(d). They are thus constrained within a smaller Hilbert space compared to chaotic states [39].
@@
II.3 Spin-wave approximation for the marked states in xorX model
The numerical results presented in the previous section suggest a more detailed study on the nature of the additional non-thermal states found by QCNN in the xorX model in Eq. (1). The PR indicates that these states are predominantly localized within a small subregion of the full Hilbert space. In this section, we demonstrate that some of these non-thermal states can be understood in terms of quasiparticles, specifically magnon bound states. We will construct effective tight-binding Hamiltonians that approximately describe the spin-wave modes of these quasiparticles, reproducing key features of the exact many-body eigenstates.
II.3.1 Integrable states
Figure 5: Domain wall dynamics between two antiferromagnetically ordered domains, with the inset highlighting the eigenstates within this subspace. These eigenstates are characterized by distinct eigen-energies with similar entanglement entropy.
@@ II.3.1 Integrable states
$H=\sum_{i=1}^{n-1}(-1)^{i}\Delta d_{i}^{\dagger}d_{i}+\lambda d_{i+1}^{\dagger}d_{i}+\lambda d_{i}^{\dagger}d_{i+1}\leavevmode\nobreak\ .$ (6)
@@
$H$ $=$ $\left(\begin{matrix}\Delta&\lambda+\lambda e^{ik}\\ \lambda+\lambda e^{-ik}&-\Delta\end{matrix}\right)\leavevmode\nobreak\ ,$ (7)
where we have set the lattice spacing to unity. The eigenenergies are $E_{k}=\pm\sqrt{\Delta^{2}+4\lambda^{2}\cos^{2}(k/2)}$, with corresponding eigenstates $\phi_{k}$. Under OBC, the eigenstates approximate standing waves, expressed as superpositions of $\phi_{k}$ and $\phi_{-k}$. Specifically, these superpositions take the form of (|$\left(\ket{\phi_{k}}+\ket{\phi_{-k}}\right)/\sqrt{2}$ and (|$\left(\ket{\phi_{k}}-\ket{\phi_{-k}}\right)/\sqrt{2}$. These states exhibit low entanglement entropy, characteristic of integrable systems, as shown in the inset of Fig. 5. The QCNN successfully identifies states within this integrable subspace, marking those with energies similar to the exact scar states.
Figure 6: The mean and variance of the total $S_{z}$ for each eigenstate within a certain $n_{\rm dw}$ sector: (a)&(c) $n_{\rm dw}=2$; (b)&(d) $n_{\rm dw}=3$. The red crosses indicate the eigenstates identified by the QCNN. The green solid curves represent analytical results from the ferromagnetic magnon bound state approximation, while the purple dashed curves correspond to those obtained from the anti-ferromagnetic magnon bound state approximation. Other parameters are $\lambda=J=10\Delta$.
II.3.2 Approximate quasiparticle states